### v2

# Dual-Expert Consistency Model for Efficient and High-Quality Video Generation

Abstract
Diffusion Models have achieved remarkable results in video synthesis but require iterative denoising steps, leading to substantial computational overhead. Consistency Models have made significant progress in accelerating diffusion models. However, directly applying them to video diffusion models often results in severe degradation of temporal consistency and appearance details. In this paper, by analyzing the training dynamics of Consistency Models, we identify a key conflicting learning dynamics during the distillation process: there is a significant discrepancy in the optimization gradients and loss contributions across different timesteps. This discrepancy prevents the distilled student model from achieving an optimal state, leading to compromised temporal consistency and degraded appearance details. To address this issue, we propose a parameter-efficient Dual-Expert Consistency Model (DCM), where a semantic expert focuses on learning semantic layout and motion, while a detail expert specializes in fine detail refinement. Furthermore, we introduce Temporal Coherence Loss to improve motion consistency for the semantic expert and apply GAN and Feature Matching Loss to enhance the synthesis quality of the detail expert. Our approach achieves state-of-the-art visual quality with significantly reduced sampling steps, demonstrating the effectiveness of expert specialization in video diffusion model distillation. Our code and models are available at https://github.com/Vchitect/DCM.
Figure 1: Comparison of visual results between our DCM (4 steps), the original HunyuanVideo, and other competing methods (left). Comparison of latency and VBench score across different methods (right). Latency is measured on two A100 GPUs under the video synthesis configuration of 129 frames at $1280\times 720$ resolution.
footnotetext: *Equal Contribution. Project Leader Corresponding Author.
1 Introduction
Diffusion Models [8] have achieved remarkable progress in image and video synthesis [48, 45, 63]. However, they require multiple iterations to model the probability flow Ordinary Differential Equation (ODE) [52] and rely on increasingly large denoising networks, resulting in substantial computational overhead that limits their practicality in real-world applications.
To mitigate this constraint, Consistency Distillation [54] has emerged as an efficient knowledge distillation framework to reduce the sampling timesteps. It leverages a pre-trained diffusion model as the teacher and trains a student model to directly map any point along the ODE trajectory to the same solution, thereby ensuring the self-consistency property. Despite enabling few-step sampling, it often struggles with visual quality, particularly in challenging video synthesis, leading to distorted layouts, unnatural motion, and degraded details.
To ground this inherent issue, we analyzed the sampling dynamics of video diffusion models, as shown in Fig. 2 (a). Our key observations are that the differences between adjacent steps are substantial in the early stages of sampling, whereas the changes become more gradual in the later stages. This discrepancy arises because the early steps primarily focus on synthesizing semantic layout and motion, while the later steps emphasize refining fine details. These findings suggest that the student model may learn different patterns and exhibit distinct learning dynamics when trained on high-noise and low-noise samples. We visualized the magnitude and gradient of the consistency loss during the distillation process and observed significant differences between high and low noise levels, as shown in Fig. 2 (b). This variation indicates that jointly distilling a single student model to capture both semantic layout and fine-detail synthesis may introduce optimization interference, potentially leading to suboptimal results.
To validate this assumption, we trained two expert denoisers. We first divide the ODE trajectory of the pre-trained model into two phases: the semantic synthesis phase and the detail refinement phase. We then train two distinct student expert denoisers, each responsible for fitting one of these sub-trajectories. During inference, we dynamically select the corresponding expert denoiser based on the noise level of samples to predict the next position in the ODE trajectory. The results demonstrate that the combination of the two student expert denoisers achieves better performance, thereby confirming the validity of our hypothesis.
However, this straightforward baseline involves training two student models which is not efficient enough. To further enhance parameter efficiency, we analyze the parameter differences between the two expert denoisers and identify that the primary differences lie in 1) embedding layers where the input parameters include timesteps, and 2) the linear layers within the attention layers. Based on this insight, we propose a parameter-efficient Dual-Expert Consistency Model (DCM). Specifically, we first train a semantic expert denoiser on the semantic synthesis trajectory. We then freeze this expert and introduce a new set of timestep-dependent layers, incorporating a LoRA [11] into the linear layers of the attention blocks. Subsequently, we fine-tune these newly added layers on the detail refinement trajectory. In this manner, we decouple the optimization of the two expert denoisers with minimal additional parameters and computational cost, achieving visual results comparable to those obtained with two separate experts.
Given the differing training dynamics of the semantic and detail expert denoisers, we introduce distinct optimization objectives beyond the original consistency loss. To enforce temporal coherence in the semantic expert denoiser, we introduce a Temporal Coherence Loss, which guides it to capture motion variations across frames. To enhance the fine-grained content synthesized by the detail expert denoiser, we introduce a generative adversarial (GAN) [4] loss and incorporate a Feature Matching loss. Specifically, we alternately optimize the student model and the discriminator in the feature space, encouraging the generator to synthesize visual content that aligns with the output distribution of the teacher model. The Feature Matching term enhances supervision over intermediate features, thereby stabilizing the GAN training.
Our proposed DCM accelerates sampling while preserving both semantic and detail quality, as shown in Fig. 1. In summary, our contributions are as follows:
  • We analyze the training dynamics of Consistency Models and identify a key conflict in the distillation process: discrepancies in loss contributions and optimization gradients across noise levels hinder optimal learning, leading to suboptimal visual quality.
  • We propose a parameter-efficient Dual-Expert Consistency Model that decouples the expert denoisers distillation, mitigating the conflict and improving visual quality with minimal parameter and computational cost.
  • To enhance visual quality, we introduce Temporal Coherence Loss for the semantic expert and GAN loss with Feature Matching term for the details expert, improving both temporal consistency and detail quality.
2 Related Work
Figure 2: Visualization of the video synthesis process and the trend of loss variation. (a) In the early stages of sampling, the results change significantly and rapidly, whereas in the later stages, the changes become gradual and smooth. (b) During distillation, the loss and gradient norm of the student model exhibit significant differences between samples with high and low noise levels.
2.1 Diffusion Models For Video Synthesis
Video Diffusion Models have witnessed rapid advancements with diffusion models [8, 9, 2], driving a wide range of visual content generation applications[70, 12, 71, 13, 72]. Building on the Diffusion Transformer (DiT) [45] pre-training, a notable breakthrough is the development of high-fidelity video diffusion models [44, 21, 55, 42, 22, 10, 63, 20]. However, scaling these models for long videos incurs significant training and inference costs. LTX-Video [7] designed Video-VAE that achieves a high compression ratio for efficient self-attention. Pyramid flow [15] introduced a unified pyramidal flow matching algorithm for efficient video generative modeling. The acceleration method based on sparse attention [68, 69] and feature cache [73, 41] has also improved inference efficiency. Moreover, while efficient fast diffusion samplers [52, 28, 32, 33, 17] reduce inference steps, further reduction often severely degrades performance. Diffusion distillation offers a promising way to reduce sampling steps while maintaining visual quality.
2.2 Diffusion Model Distillation
Diffusion distillation [37] aims to distill knowledge from pre-trained diffusion models to student models, reducing inference cost. Prior works can be generally classified into two categories based on their distillation mechanisms.
Trajectory-preserving distillation methods exploit the fact that diffusion models learn an Ordinary Differential Equation (ODE) trajectory and aim to predict the exact teacher output in fewer steps. Among the earliest studies on diffusion distillation, Luhman et al. [34] and DSNO [74] proposed training the student model using noise-image pairs precomputed by the teacher model with an ODE solver. Progressive distillation [49, 43] reduces the final number of sampling steps by iteratively applying the distillation process to halve the number of sampling steps of previous model. Instaflow [29, 30] progressively learns straighter flows, enabling accurate one-step predictions over larger distances. Consistency models [54, 53, 35, 36, 27, 75, 31], BOOT [5] and TRACT [1] learn to map samples along the ODE trajectory to another point to achieve self-consistency. The consistency trajectory model [18] was designed to mitigate discretization inaccuracies and accumulated estimation errors in the multistep consistency model sampling. PCMs [57] phase the ODE trajectory into several sub-trajectories and only enforce the self-consistency property on each sub-trajectory, thus alleviating the limitations of CMs. TCM [23] generalizes consistency training to the truncated time range to prevents the truncated-time training from collapsing to a trivial solution. Trajectory-preserving distillation enables stable optimization but can degrade visual quality, leading to blurriness or distortions when sampling with fewer steps.
Distribution-matching distillation methods bypass the ODE trajectory and aim to train the student model to generate samples whose distribution aligns with that of the teacher diffusion model. Some methods reduce the distribution gap between the student and teacher models through adversarial training [61, 40, 16, 51, 50, 26, 3, 62]. Other methods [38, 39, 65, 64, 66, 76] achieve diffusion distillation by score distillation [46]. Notably, DMD [65] aligns the one-step generator with the distribution of teacher model by minimizing an approximate KL divergence, whose gradient is the difference between the target and synthetic distribution score functions. Recently works [25, 47, 19] have also tried to integrate the advantages of trajectory-preserving and distribution-matching methods. Hyper-SD [47] introduced the trajectory-segmented consistency distillation and used DMD [65] for one-step generation enhancement.
Previous works have primarily focused on distilling image synthesis diffusion models, with some efforts [24, 58, 67] extending to the distillation of small-scale video synthesis models [6, 59]. However, these methods are limited to synthesizing low-resolution and short-sequence videos. Seaweed-APT [26] proposed adversarial post-tuning against real data following diffusion pre-training for one-step high-resolution 2-second duration video generation. A recent work [66] extended DMD [65] for video synthesis in four sampling steps. Due to the inherent complexity of video synthesis and the increasing model scale, research on diffusion distillation for video synthesis remains limited, and its performance is yet to be fully explored.
3 Methodology
Figure 3: Comparison of the visual quality of denoiser variants trained at different noise level samples. By optimizing two expert denoisers to decouple the distillation process into semantic learning and detail learning, and combining them during inference, we achieve the best quantitative and qualitative visual results.
3.1 Preliminary
Diffusion Model is a generative framework with a forward and reverse process.
In the forward process, noise is progressively added to clean data $\boldsymbol{x}_{0}\sim p_{\mathrm{data}}(\boldsymbol{x}_{0})$, degrading the signal:
$q(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})=\mathcal{N}(\boldsymbol{x}_{t};\sqrt{\alpha_{t}}\boldsymbol{x}_{0},\sqrt{1-\alpha_{t}}\boldsymbol{I}),$ (1)
where $\{\alpha_{t}\}_{t=1}^{T}$ controls the noise schedule.
The reverse process, typically parameterized by a UNet or transformer $\boldsymbol{\epsilon}_{\theta}$, is trained to predict the noise:
[…]
During inference, a clean sample $\boldsymbol{x}_{0}$ can be recovered through iterative denoising:
$p(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t})=\mathcal{N}(\boldsymbol{x}_{t-1};\mu_{\theta}(\boldsymbol{x}_{t},t),\Sigma_{\theta}(\boldsymbol{x}_{t},t)),$ (3)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are learned parameters.
Consistency Distillation utilizes a pre-trained model $\boldsymbol{\epsilon}_{\theta}$ as the teacher $F_{T}$ to distill its knowledge into a student model $F_{S}$ initialized with $\boldsymbol{\epsilon}_{\theta}$, allowing for faster sampling with fewer steps [54]. Specifically, consistency distillation trains the student model $F_{S}$ to directly map any point $\boldsymbol{x}_{t_{n}}$ on the solution trajectory of the ODE solver $\Phi$ to its endpoint $\boldsymbol{x}_{t_{end}}$. The learning objective can be formulated as:
$\mathcal{L}_{CD}=\mathbb{E}_{\boldsymbol{x},t_{n}}||\Phi(\boldsymbol{x}_{t_{n}},F_{S}(\boldsymbol{x}_{t_{n}},t_{n},c),t_{end})$- (4)
$\Phi(\hat{\boldsymbol{x}}_{t_{n-1}},F_{S}^{-}(\hat{\boldsymbol{x}}_{t_{n-1}},t_{n-1},c),t_{end})||_{2}^{2}.$
Here, $F_{S}^{-}$ is the exponential moving average (EMA) of $F_{S}$ and $\hat{\boldsymbol{x}}_{t_{n-1}}$ is the next point on the ODE solution trajectory computed by the teacher model $F_{T}$:
$\hat{\boldsymbol{x}}_{t_{n-1}}=\Phi(\boldsymbol{x}_{t_{n}},F_{T}(\boldsymbol{x}_{t_{n}},t_{n},c),t_{n-1}).$ (5)
Consistency distillation has garnered widespread attention and research [35, 47, 57] due to its ease of stable training.
Figure 4: The training process of DCM consists of two stages. In the semantic learning stage, we train SemE on high-noise samples with consistency loss and temporal coherence loss as the learning objectives. In the detail learning stage, we initialize DetE with the weights of SemE and introduce a set of time-dependent layers and LoRA. DetE is then trained on low-noise samples, where only the newly added layers and LoRA are updated. The learning objectives in this stage include consistency loss, GAN loss, and Feature Matching loss.
Figure 5: Weight difference distribution between expert denoisers. We employ the normalized L1 distance to quantify the difference between the weights.
3.2 Suboptimal Solution in Consistency Distillation
Although consistency distillation has demonstrated promising results in class-conditioned image synthesis and text-to-image model distillation, it falls short in more challenging video synthesis diffusion models, where issues arise such as distorted layouts, unnatural motion, and degraded details.
Due to the limited capacity of the student model, consistency distillation struggles to address these issues simultaneously. By tracking the video synthesis process, we find that in the early stages of sampling, the sampling results vary significantly and rapidly, whereas in the later stages, the transitions become more gradual and smooth, as shown in Fig. 2 (a). In the early stages of sampling, rapid changes establish the semantic layout and motion of the video, while in the later stages, the model gradually refines the details with smooth adjustments. These findings imply that the student model may learn different patterns with distinct training dynamics when distilled with high-noise and low-noise samples. By visualizing the trends of consistency loss and gradient norm during distillation, we find significant differences in the student model when distilled with high- and low-noise samples, as shown in Fig. 2 (b). This suggests that jointly optimizing a student model for both semantic and fine-grained details synthesis may introduce inefficient optimization, constraining its fitting capacity and leading to suboptimal performance.
To validate our hypothesis, we conducted an experiment on the HunyuanVideo [20] text-to-video diffusion model. Specifically, we divided the ODE solution trajectory ($\boldsymbol{x}_{N},\boldsymbol{x}_{N-1},...,\boldsymbol{x}_{1},\boldsymbol{x}_{0}$) of the pre-trained model into two sub-trajectories, using $t_{\kappa}$ as the boundary (we set $N=50$ and $\kappa=37$ by default). The first part ($\{\boldsymbol{x}_{t_{i}}\}_{i=\kappa}^{N}$) primarily focuses on synthesizing the semantic layout and motion, while the second part ($\{\boldsymbol{x}_{t_{j}}\}_{j=0}^{\kappa}$) emphasizes semantic refinement and high-quality detail generation. As shown in Fig 3 (a), we optimized two distinct student models, semantic expert denoiser $F_{\text{SemE}}$ and details expert denoiser $F_{\text{DetE}}$, to fit each sub-trajectory:
$\mathcal{L}_{\text{SemE}}=\mathbb{E}_{\boldsymbol{x},t_{m}\in[t_{\kappa},t_{N}]}||\Phi(\boldsymbol{x}_{t_{m}},F_{\text{SemE}}(\boldsymbol{x}_{t_{m}},t_{m},c),t_{\kappa})$- (6)
$\Phi(\hat{\boldsymbol{x}}_{t_{m-1}},F_{\text{SemE}}^{-}(\hat{\boldsymbol{x}}_{t_{m-1}},t_{m-1},c),t_{\kappa})||_{2}^{2},$
$\mathcal{L}_{\text{DetE}}=\mathbb{E}_{\boldsymbol{x},t_{n}\in[t_{0},t_{\kappa}]}||\Phi(\boldsymbol{x}_{t_{n}},F_{\text{DetE}}(\boldsymbol{x}_{t_{n}},t_{n},c),t_{0})$- (7)
$\Phi(\hat{\boldsymbol{x}}_{t_{n-1}},F_{\text{DetE}}^{-}(\hat{\boldsymbol{x}}_{t_{n-1}},t_{n-1},c),t_{0})||_{2}^{2}.$
Expert denoiser $F_{\text{SemE}}$ is optimized to synthesize coherent semantic layouts and motion, while expert denoiser $F_{\text{DetE}}$ learns to generate high-quality details. During inference, we dynamically switch the expert denoiser based on the sampling stage. To assess the impact of each expert denoiser and the effectiveness of decoupled training, we evaluate four variants: a) VCM: The vanilla consistency model is used throughout the sampling process. b) SemE + VCM: SemE is applied in the first sub-trajectory, transitioning to VCM in the second. c) VCM + DetE: VCM is applied initially, followed by DetE. d) SemE + DetE: SemE and DetE are integrated for the sampling process. According to the Fig. 3 (b) and (c), SemE and DetE have respectively learned to model semantics and details, each outperforming VCM in their respective aspects. This validates our hypothesis, demonstrating that decoupled optimization is superior to jointly training a single model for both tasks.
3.3 Parameter-efficient Dual-Expert Distillation
While training two expert denoisers improves video quality, it significantly increases model parameters and GPU memory consumption during inference. Through the analysis of parameter similarity between the two expert denoisers, we found that the primary differences in model parameters lie in the embedding layers $\Psi$ where the input parameters include timesteps and the linear layers within the attention layers $\Lambda$, as illustrated in Fig. 5.
Based on the above observations, we propose the parameter-efficient Dual-Expert distillation strategy, as illustrated in Fig. 4. Specifically, our training scheme is divided into two stages. 1) Initialize the semantic expert denoiser $F_{\text{SemE}}$ with the teacher model $F_{T}$ and optimize all its parameters on the sub-trajectory $\{\boldsymbol{x}_{t_{i}}\}_{i=\kappa}^{N}$. 2) Use the optimized $F_{\text{SemE}}$ as the initialization of $F_{\text{DetE}}$ and freeze it. Then add a new set of timestep-dependent embedding layers $\Psi$ and LoRA [11] $\Lambda^{\dagger}$ of attention blocks. Optimize the newly added parameters ($\Psi$ and $\Lambda^{\dagger}$) on its sub-trajectory $\{\boldsymbol{x}_{t_{j}}\}_{j=0}^{\kappa}$. In this way, we significantly reduce the number of parameters required for decoupling the optimization process of semantic modeling and detail learning, with minimal computational cost, while maintaining the visual quality of the synthesized videos.
3.4 Expert-specific Optimization Objective
In addition to the consistency objectives mentioned in Eq. 6 and Eq. 7, we also designed expert-specific optimization objectives for the semantic expert denoiser $F_{\text{SemE}}$ and details expert denoiser $F_{\text{DetE}}$, as shown in Fig. 4.
Temporal Coherence Loss To enhance the temporal coherence in the video synthesized by the semantic expert denoiser $F_{\text{SemE}}$, we introduce the Temporal Coherence Loss $\mathcal{L}_{TC}$, which emphasizes and guides the $F_{\text{SemE}}$ to focus on the variations and motion at corresponding positions among different frames:
$\boldsymbol{x}_{t_{\kappa}}$ $=\Phi(\boldsymbol{x}_{t_{m}},F_{\text{SemE}}(\boldsymbol{x}_{t_{m}},t_{m},c),t_{\kappa}),$ (8)
$\hat{\boldsymbol{x}}_{t_{\kappa}}$ $=\Phi(\hat{\boldsymbol{x}}_{t_{m-1}},F_{\text{SemE}}^{-}(\hat{\boldsymbol{x}}_{t_{m-1}},t_{m-1},c),t_{\kappa}),$
[…]
Here $\boldsymbol{x}^{t_{\kappa}}_{l:L}$ represents the video latents from the $l$-th to the $L$-th channel along the temporal axis. This temporal coherence loss encourages the semantic expert denoiser $F_{\text{SemE}}$ to preserve consistent motion and spatial relationships between frames, ensuring more fluid videos synthesis.
Generative Adversarial Loss The effectiveness of the generative adversarial (GAN) [4] loss in high-quality detail synthesis has been validated in many distribution-matching distillation methods. We introduce the GAN loss into the training of the details expert denoiser and incorporate a Feature Matching loss to stabilize the training. We first obtain $\boldsymbol{x}_{t_{0}}$ and $\hat{\boldsymbol{x}}_{t_{0}}$ with the details expert denoiser $F_{\text{DetE}}$, teacher model $F_{T}$ and ODE solver $\Phi$:
$\boldsymbol{x}_{t_{0}}$ $=\Phi(\boldsymbol{x}_{t_{n}},F_{\text{DetE}}(\boldsymbol{x}_{t_{n}},t_{n},c),t_{0}),$ (9)
$\hat{\boldsymbol{x}}_{t_{0}}$ $=\Phi(\hat{\boldsymbol{x}}_{t_{n-1}},F_{\text{DetE}}^{-}(\hat{\boldsymbol{x}}_{t_{n-1}},t_{n-1},c),t_{0}),$
Then we perform the forward process and apply noise to them to obtain fake sample $\boldsymbol{x}_{fake}$ and real sample $\boldsymbol{x}_{real}$ with Eq. 1. We use a frozen teacher model as the feature extraction backbone $\Omega$, extracting intermediate features with a fixed stride for calculating the GAN loss and Feature Matching loss $\mathcal{L}_{FM}$. During training, we iteratively update the parameters of $F_{\text{DetE}}$ and the discriminator head $f_{D}$:
$\mathcal{L}_{FM}=\mathbb{E}_{\boldsymbol{x},t_{n}}\left\|\Omega(\boldsymbol{x}_{fake})-\Omega(\boldsymbol{x}_{real})\right\|^{2}_{2},$ (10)
$\mathcal{L}_{G}=\mathbb{E}_{\boldsymbol{x},t_{n}}[1-f_{D}(\Omega(\boldsymbol{x}_{fake}))]+\mathcal{L}_{FM},$
$\mathcal{L}_{D}=\mathbb{E}_{\boldsymbol{x},t_{n}}[f_{D}(\Omega(\boldsymbol{x}_{fake}))]+\mathbb{E}_{\boldsymbol{x},t_{n}}[1-f_{D}(\Omega(\boldsymbol{x}_{real}))].$
The integration of the GAN loss in combination with Feature Matching loss provides a robust framework for training the details expert denoiser $F_{\text{DetE}}$, stabilizing its learning process and improving the quality of detail synthesis.
4 Experiments
4.1 Experimental Setup
Backbones and Baselines We utilize HunyuanVideo [20] and CogVideoX [63] as the base models for distillation. The HunyuanVideo has 13 billion parameters, and CogVideoX has 2 billion parameters. Since most prior distillation methods for diffusion models have not been applied to video synthesis, we follow the official implementations of LCM [35] and PCM [57] to implement these two methods on the selected base models as baselines for comparison.
Implementation Details For HunyuanVideo, we selected trajectories with 50 Euler steps and used the default sampling parameters from diffusers. The distillation was conducted on 129-frame video sequences at a resolution of $1280\times 720$ with a batch size of 6. For the semantic expert denoiser, we performed $1000$ iterations of distillation with a learning rate of 1e61e-6, while for the details expert denoiser, we trained for $1000$ iterations with a learning rate of 5e65e-6. For CogVideoX-2B, we selected trajectories with 50 DDIM steps. The distillation was conducted on 29-frame video sequences at a resolution of $720\times 480$ with a batch size of 4. In the first-stage fine-tuning, we distilled for approximately 1000 steps, while in the second-stage fine-tuning, we distilled for around 500 steps, both with a learning rate of 1e61e-6. All experiments were conducted on 24 NVIDIA A100 80GB GPUs.
Evaluation Metrics For video quality evaluation, we use VBench [14] as our assessment metric. VBench is a comprehensive benchmark suite for video generative models, designed to align closely with human perception and offer valuable insights from multiple perspectives. Additionally, we conducted a user study to help evaluate the visual quality of the generated videos.
4.2 Main Results
Quantitative Comparison Table 1 presents the quantitative comparison of our method with LCM and PCM on HunyuanVideo and CogVideoX. We generate videos using the prompts provided by VBench to evaluate their performance in terms of semantic alignment and visual quality. It can be observed that on HunyuanVideo, our method achieves a VBench score comparable to the baseline with 4-step sampling, significantly outperforming LCM and PCM. In terms of efficiency, our method incurs a nearly identical latency cost per inference step compared to LCM and PCM.
Figure 6: Visual quality comparison of different methods. Differences are highlighted in boxes.
Table 1: Comparison of efficiency and visual quality of different methods. The latency of HunyuanVideo was measured on two A100 GPUs, and that of CogVideoX on a single A100 GPU.
Method Step Lat.(Sec.) VBench
Total Quality Semantic
[…]
LCM 4 120.68 80.33 80.83 78.32
PCM 4 120.89 80.93 81.94 76.90
Ours 4 121.52 83.83 85.12 78.67
LCM 8 242.80 81.49 82.35 78.03
PCM 8 242.96 81.63 82.78 77.00
Ours 8 244.72 83.86 85.00 79.32
[…]
LCM 4 3.22 78.88 80.07 74.12
PCM 4 3.23 79.09 80.33 74.14
Ours 4 3.31 79.99 81.35 74.56
LCM 8 6.42 79.34 80.64 74.21
PCM 8 6.42 79.70 80.98 74.60
Ours 8 6.58 80.26 81.57 75.03
Qualitative Comparison Fig. 6 presents a comparison of the videos generated by our method and those produced by the original model, LCM and PCM. The results demonstrate that our method maintains high semantic and detail quality in synthesized videos while reducing the number of inference steps. Additional qualitative results are provided in the supplementary material for further reference.
User Study To further evaluate the effectiveness of our method, we conduct a human evaluation to assess the perceived visual quality of the generated videos. Specifically, we randomly select 30 videos for each model. During the evaluation, each rater is presented with a text prompt along with two videos generated by different distillation methods, displayed in a randomized order to eliminate bias. Following the protocol of human preference evaluation in HunyuanVideo [20], the professional raters are asked to choose the video they perceive to have superior text alignment, motion quality, and visual quality. Each sample is evaluated by fifty independent raters, and the aggregated voting results are summarized in Table 2. As one can see, compared to other distillation methods, the raters significantly prefer the videos generated by our method.
Table 2: User preference study. The numbers represent the percentage of raters who favor the videos synthesized by our method.
Method comparison HunyuanVideo CogVideoX
Ours vs. LCM 82.67% 75.33%
Ours vs. PCM 77.33% 72.67%
4.3 Ablation Study
To thoroughly evaluate both the effectiveness of our method, we conduct extensive ablation studies based on HunyuanVideo, as shown in Table 3. All experiments were conducted on 29-frame videos with a resolution of $1280\times 720$. Inference is performed with 4 sampling timesteps.
Table 3: Impact of different components of our method.
Variants VBench
OD PE TC GF Total Quality Semantic
(1) 80.30 80.74 78.36
(2) 83.08 84.20 78.59
(3) 83.03 84.16 78.53
(4) 83.42 84.63 78.63
(5) 83.71 84.99 78.59
(6) 83.80 85.10 78.62
Figure 7: Impact of optimization decoupling and parameter-efficient distillation.
Effect of Optimization Decoupling (OD) Through Experiments (1) and (2), decoupling the optimization of semantic and detail modeling significantly improves the semantic and quality scores. As shown in Fig. 7, the optimized decoupled model synthesizes videos with better semantic and detail quality, where the motion of characters and facial details appear more natural.
Effect of Parameter-Efficient dual-expert distillation (PE) Through Experiments (2) and (3), we observe that compared to simply decoupling the optimization into two separate model training processes, the parameter-efficient Dual-Expert distillation significantly reduces both the parameters and memory requirements, with minimal computational overhead, while preserving visual quality. The last two rows of Fig. 7 also demonstrate that our parameter-efficient Dual-Expert method does not result in a significant degradation in visual quality.
Figure 8: Impact of temporal coherence loss.
Effect of Temporal Coherence Loss (TC) By comparing Experiments (3) and (4), or (5) and (6), we observe that the introduction of Temporal Coherence loss improves the quality scores of the synthesized videos. As shown in Fig. 8, the introduction of the TC loss enables more natural motion in the video and enhances its consistency.
Figure 9: Impact of the GAN loss and Feature Matching term.
Effect of GAN and Feature Matching Loss (GF) By comparing Experiments (3) and (5), or (4) and (6), we observe that the introduction of GAN Loss improves the quality scores of the synthesized videos. As shown in Fig. 9, the introduction of the GAN loss and Feature Matching term enhances the realism of the details in the synthesized video.
Figure 10: Impact of different $\kappa$.
Selection of $\kappa$ In this paper, we determine the value of $\kappa$ based on the inference process. Fig. 10 (left) illustrates the L1 distance between adjacent time-step sampling results in HunyuanVideo during the sampling process. It can be observed that from approximately step 37 onward, the L1 Distance decreases to a very small value. We interpret this as the point where the semantic content and layout are established, and the remaining steps focus on synthesizing high-frequency details. Therefore, we set $\kappa=37$ as the default value. To evaluate the impact of different $\kappa$ values, we experimented with $\kappa=28,35,37,39,46$. As shown in Fig. 10 (right), the results indicate that as $\kappa$ deviates from the transition point between semantic synthesis and detail synthesis, the video quality gradually deteriorates. It further validates the effectiveness of our optimization decoupling strategy.
5 Conclusion and Discussion
In this paper, we identify a key optimization conflict in consistency distillation for video synthesis: there exists a significant discrepancy in the optimization gradients and loss contributions across different timesteps. Distilling the entire ODE trajectory into a single student model fails to balance these aspects, leading to degraded motion consistency and coarse synthesis quality. To address this issue, we propose a parameter-efficient Dual-Expert distillation framework that decouples semantic learning from fine-detail refinement. Additionally, we introduce a Temporal Coherence loss to enhance motion consistency for the semantic expert and apply GAN and Feature Matching loss to improve synthesis quality for the detail expert. Our method significantly reduces sampling steps while achieving state-of-the-art visual quality, demonstrating the effectiveness of expert specialization in video diffusion model distillation.
Limitation Although our method achieves favorable results with 4 steps inference, it still struggles to produce satisfactory outcomes with fewer steps (e.g., 2) due to limited training data and iterations. We will further explore high-quality synthesis with fewer steps in future work.
6 Acknowledgement
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Supplementary Material
7 Further implementation details
Stage division and expert switching. During inference, we empirically observe that evenly dividing the total steps between the two experts produces favorable results. With 8 or 4 total steps, we assign 4 or 2 steps to each expert, respectively. These steps are uniformly sampled within each sub-trajectory.
8 Additional Results
8.1 Compatibility with other acceleration techniques
DCM accelerates generation via sampling step reduction and is compatible with other methods like low precision computation and sparse modeling. For example, integrating SVG [60] (which leverages the sparsity of 3D full attention), yields an additional 1.33× speedup on top of DCM-Hunyuan while maintaining high fidelity (VBench 83.79%).
8.2 Generality of DCM
DCM addresses discrepancies in loss and gradient contributions across noise levels—a problem inherent to consistency distillation itself, not from any specific model architecture. Beyond HunyuanVideo [20] and CogVideoX [63], we further apply it to the recent WAN2.1-T2V [56]. DCM significantly accelerates inference while preserving comparable visual quality, as evidenced by VBench scores (baseline: 83.2%, DCM: 82.9%).
8.3 Visualization of the sampling process
To further verify the effectiveness of our method in semantic and detail synthesis, we visualize the results of each sampling step in a 4-step sampling process on HunyuanVideo. As shown in Fig. 11 and Fig. 12, our method achieves better performance in both semantic layout and fine details compared to competing methods.
Figure 11: Visualization of the sampling process of different methods.
Figure 12: Visualization of the sampling process of different methods.
8.4 More visual comparison results
The additional visual comparison results for HunyuanVideo are presented in Fig. 13, Fig. 14 and Fig. 15. More visual results of CogVideoX are shown in Fig. 16. The results indicate that our method maintains reliable fidelity across diverse models, styles, and content in video synthesis while also achieving acceleration.
Figure 13: Visual quality comparison of different methods.
Figure 14: Visual quality comparison of different methods.
Figure 15: Visual quality comparison of different methods.
Figure 16: Visual quality comparison of different methods.